A particular algorithm increases in time as the number of operations, n, increases.


Suppose the time complexity of this algorithm is given by: f(n) =4n3+5n2 * log(n).

Show that f(n) is O(g(n)) for g(n) = n3.



1
Expert's answer
2021-10-25T17:27:35-0400

"f(x)=O(g(x))" if there exists a positive real number M and a real number x0 such that

"|f(x)|\le Cg(x)" for all "x\ge x_0"


So, we have:

f(n) =4n3+5n2 * log(n) < "5g(x)=5n^3" for "x\ge 2"

where "C=5,x_0=2"